OpenStudy (anonymous):

What is (x^2+1) times (2x)

OpenStudy (unklerhaukus):

distribute (a+b) x (c) = (a x c) + (b x c)

OpenStudy (anonymous):

So (x^2 (times) 2x) + (1 (times) 2x)

OpenStudy (unklerhaukus):

That's right

OpenStudy (anonymous):

Well I don't know what the first half would be. I don't know how to do the x's So it's something + 2x

OpenStudy (unklerhaukus):

ok well your almost there

OpenStudy (unklerhaukus):

\[An^p\times Bn^q=(A\times B)\times n^{p+q} \]

OpenStudy (anonymous):

2x^3?

OpenStudy (unklerhaukus):

YES!

OpenStudy (anonymous):

Well that's what I thought it was but it's not one of my choices...

OpenStudy (unklerhaukus):

what was your final answer?

OpenStudy (anonymous):

2x^3 + 2x

OpenStudy (unklerhaukus):

Hmm, that is right, what are the choices?

OpenStudy (anonymous):

2x^2 +2 2x^2+1 x^2+4x+4 4x^2+1

OpenStudy (unklerhaukus):

Are you sure those are the right options for this question?

OpenStudy (anonymous):

Yepp. Here I'll put the whole question here so you can see if I'm trying to do this wrong.

OpenStudy (unklerhaukus):

ah i see the propbem

OpenStudy (unklerhaukus):

you have found the product of the functions \[(f\cdot g)(x)=f(x)\times g(x)\] the question is looking for the function of a fcunction \[[f\circ g](x)=f\big(g(x)\big)\]

OpenStudy (unklerhaukus):

*function of a function.

OpenStudy (unklerhaukus):

\[f(x)=x^2+1\\g(x)=2x\\ \\ [f\circ g](x)=f\big(g(x)\big)=f(2x)=(2x)^2+1=\]

OpenStudy (unklerhaukus):

\[[g\circ g](x)=g\big(g(x)\big)=g(2x)=2(2x)=4x\]

OpenStudy (unklerhaukus):

have you seen this notation before?

OpenStudy (unklerhaukus):

@RoseDryer

OpenStudy (unklerhaukus):

are you still here?

OpenStudy (anonymous):

I'm sorry I was making a bottle.

OpenStudy (anonymous):

I have not seen that before

OpenStudy (unklerhaukus):

does it make any sense ?

OpenStudy (anonymous):

So it would be f(g(x))

OpenStudy (unklerhaukus):

yeah

OpenStudy (anonymous):

So (2x)^2+1

OpenStudy (unklerhaukus):

yes, last step is to simplify that

OpenStudy (anonymous):

How do I?

OpenStudy (unklerhaukus):

\[(2x)^2=(2x)\times(2x)=(2\times2)(x\times x)=\]

OpenStudy (anonymous):

So 4x^2?

OpenStudy (unklerhaukus):

yep

OpenStudy (unklerhaukus):

so your final answer now is ____ _ _

OpenStudy (unklerhaukus):

yep

OpenStudy (anonymous):

4x^2+1

OpenStudy (unklerhaukus):

Correct \[\boxed{\large\color{red}\checkmark}%unk\]

OpenStudy (unklerhaukus):

do you want another simple example with this new notation?

OpenStudy (anonymous):

I just want to know why we simplified.

OpenStudy (unklerhaukus):

do you when we simplified (2x)^2 ?

OpenStudy (anonymous):

What?

OpenStudy (unklerhaukus):

do you mean* when we simplified (2x)^2 ?

OpenStudy (anonymous):

Yeah. Why did we have to do that?

OpenStudy (unklerhaukus):

we did that because our final answer is a quadratic, and a nice general for of a quadratic is ax^2+bx+c=0 where a, b, c are the coefficeints, we simplified to get the coefficent next to the x^2 term

OpenStudy (anonymous):

Ahh. Okay

OpenStudy (unklerhaukus):

if (2x)^2+1 was an option that would be correct as well.

OpenStudy (anonymous):

It is.

OpenStudy (unklerhaukus):

\(2x^2\) is different to \((2x)^2\)

OpenStudy (anonymous):

Oh okay.